Sandro invested Php38,000 in the International Bank of the Philippines at 5% interest. How much additional money must he invest at 10% so that the annual income would be equal to 7.5% of his investment?
Answer by lwsshak3(11628)

Sandro invested Php38,000 in the International Bank of the Philippines at 5% interest. How much additional money must he invest at 10% so that the annual income would be equal to 7.5% of his investment?
**
let x=additional money invested at 10%
..
5%(38000)+10%x=7.5%(38000+x)
1900+0.10x=2850+.075x
.10x-.075x=2850-1900
.025x=950
x=38000
additional money invested at 10%=Php38,000